# Can Large Language Models Identify And Reason About Security Vulnerabilities? Not Yet

### v1

@@ Abstract
Large Language Models (LLMs) have been suggested for use in automated vulnerability repair, but benchmarks showing they can consistently identify security-related bugs are lacking. We thus perform the most detailed investigation to date on whether LLMs can reliably identify security-related bugs. We construct a series of 228 code scenarios and analyze eight of the most capable LLMs across eight different investigative dimensions in an automated framework. Our evaluation shows LLMs provide non-deterministic responses, incorrect and unfaithful reasoning, and perform poorly in real-world scenarios outside their knowledge cut-off date. Most importantly, our findings reveal significant non-robustness in even the most advanced models like ‘PaLM2’ and ‘GPT-4’: by merely changing function or variable names, or by the addition of library functions in the source code, these models can yield incorrect answers in and of cases, respectively. These findings demonstrate that further LLM advances are needed before LLMs can be used as general purpose security assistants.
@@ 1 Introduction
Large Language Models (LLMs) like OpenAI’s Codex [1] and GitHub Copilot [2] showed that LLMs could be used with remarkable performance for programming-language related tasks such as code generation, code documentation, and debugging. In 2022, around 1.2 million developers used Copilot, and since then we have witnessed the release of increasingly capable LLM models at a quick pace [3, 4, 5, 6]. LLMs could be particularly useful to help developers with their cybersecurity needs, as humans typically produce and miss many security relevant bugs. This issue was highlighted in the 2022 GitLab Survey [7], noting that “developers do not find enough bugs early enough” and “do not prioritize the bug remediation” when developing. It is pertinent then to investigate if LLMs could be an aid towards early identification of security problems, especially as LLMs have been already suggested for use in automated bug repair [8].
@@
In this paper, we aim to answer the following question: Can LLMs be used as helpful security assistants for vulnerability detection? This is an important question, especially as LLMs are not infallible in security-related tasks, for example introducing vulnerabilities into source code [9, 10] and software testing [11]. Unfortunately, there is no standardized and automated approach to evaluate the performance of LLMs at identifying vulnerable code. We fill this gap by introducing a generalized, fully automated, and scalable framework to systematically evaluate the performance (i.e., accuracy and reasoning capabilities) of LLMs for vulnerability detection. Our framework tests the capabilities of a given LLM as a security assistant across eight distinct dimensions: (1) deterministic response, (2) performance over range of parameters, (3) diversity of prompts, (4) faithful reasoning, (5) evaluation over variety of vulnerabilities, (6) assessment of various code difficulty levels, (7) robustness to code augmentations, and (8) use in real-world projects.
@@
We apply our framework to eight of the most capable LLMs across 228 code scenarios spanning over 8 most critical vulnerabilities in C and Python, and show that: (a) LLM performance varies widely depending on the model and the prompting technique used, however all models analyzed have a high false positives rate (FPR), and flag code where vulnerabilities have been patched as still vulnerable. (b) the output of LLMs is non-deterministic, with all models changing their answers over multiple runs for one or more of our tests. (c) even when they correctly identify a vulnerability, the reasoning that LLMs provide for this decision is often incorrect, questioning their trustworthiness. (d) LLM chain-of-thought reasoning [12] is not robust, and can be ‘confused’ by even simple code augmentations such as whitespace modification, changing function names, or using different but related library functions. (e) LLMs fail at detecting vulnerabilities in real-world scenarios that they have not seen during their training. Our study provides significant evidence that LLMs are not yet ready to be used for automated vulnerability detection, and the successful usage of our framework as a benchmark suite by future models would demonstrate meaningful progress in this space.
@@
We develop a comprehensive framework to test LLMs for their ability to identify software vulnerabilities. This framework is fully automated and includes a set of 228 code scenarios, and 17 prompting techniques. We will publicly release our framework, allowing the community to test newly developed LLMs and easily keep track of their progress in being able to identify vulnerabilities.
@@ 2 Background and Related Work
| gpt-4 |
GPT-4 |
1.76T |
8,192 |
Rem |
09/2021 |
| gpt-3.5-turbo-16k |
GPT-3.5 |
175B |
16,385 |
Rem |
09/2021 |
| codechat-bison@001 |
PaLM2 |
340B |
6,144 |
Rem |
mid-2021 |
| chat-bison@001 |
PaLM2 |
340B |
6,144 |
Rem |
mid-2021 |
| codellama-7b-instruct |
Llama2 |
7B |
100k |
Loc |
01/2022 |
| codellama-13b-instruct |
Llama2 |
13B |
100k |
Loc |
01/2022 |
| codellama-34b-instruct |
Llama2 |
34B |
100k |
Loc |
01/2022 |
| starchat-beta |
StarCoder+ |
15.5B |
8,192 |
Loc |
09/2022 |
@@
| Vuln. Code Gen. [9] |
89 |
| Vuln. Code Repair [8] |
syn. |
| Limits of ML [31] |
res. |
| Transf. Vuln. Det. [35] |
res. |
| Our Study |
228 |
Recent work has evaluated LLMs for vulnerable code generation [9], code repair [8], and vulnerability detection [35]. However, these approaches are either limited by the number of LLMs, coverage of vulnerabilities, or robustness testing. Most importantly, these studies only evaluate binary responses, checking whether the LLM gives the right label to code snippets (e.g., ‘vulnerable’ or ‘not vulnerable’). In this paper, we present the first comprehensive evaluation framework to evaluate LLMs on the task of vulnerability detection, providing a multi-faceted analysis of the capabilities of LLMs, including going beyond binary decisions and evaluating their reasoning abilities. Table II summarizes evaluation studies of LLMs, highlights their short-comings, and compares against our framework.
3 Evaluation Framework
Figure 2 presents an overview of our fully automated framework, which is designed to be applicable to any chat-based LLM. Section 3.1 describes how users can configure an LLM for integration. The core of our framework consists of five pre-defined key components: (i) a set of parameters (Section 3.2), (ii) an extensive set of prompt templates (Section 3.3), (iii) datasets (Section 3.4), and (iv) code augmentations (Section V), all of which facilitate the generation of test prompts. Each prompt is then passed to the LLM to generate a response, and the quality of the response is then analyzed by the (v) ‘Evaluator’ module (Section 3.6).
@@ 3.1 LLM configuration
To scale our evaluation framework for any chat-based LLM, we have established three configurable user inputs:
@@ 3.3 Prompt Templates
| S1 |
ZS-TO |
Code snippet is added to the input prompt with a question about a specific Common Weakness Enumeration (CWE) (e.g., out-of-bound write, path traversal). |
| S2 |
ZS-RO |
Same as S1, but the LLM is assigned the role of a ‘helpful assistant’. |
| S3 |
ZS-RO |
Similar to S1, with the LLM acting as a ‘security expert’. |
| S4 |
ZS-RO |
The LLM is defined as a ‘security expert’ who analyzes a specified security vulnerability, without the question being added to the input prompt. |
| S5 |
FS-TO |
Similar to S1, but includes a vulnerable example, its patch, and standard reasoning from the same CWE. |
| S6 |
FS-RO |
Like S4, but also includes a vulnerable example, its patch, and standard reasoning from the same CWE. |
| R1 |
ZS-TO |
Similar to S1, but begins with ”Lets think step by step” [38] to encourage a methodical approach. |
| R2 |
ZS-RO |
The LLM plays the role of a security expert with a multi-step approach to vulnerability detection, following a chain-of-thought reasoning. |
| R3 |
ZS-TO |
A multi-round conversation with the LLM, starting with a code snippet and progressively analyzing sub-components for a security vulnerability like human security-experts. |
| R4 |
FS-RO |
Similar to S6, but the reasoning for answers involves step-by-step analysis developed by the first author. |
| R5 |
FS-RO |
Like R2, but includes few-shot examples (from the same CWE) with step-by-step reasoning for detecting vulnerabilities. |
| R6 |
FS-TO |
Similar to R5, but does not assign a specific role to the LLM in the system prompt. |
| D1 |
ZS-TO |
Adds the definition of a security vulnerability to the input prompt, followed by a related question. |
| D2 |
ZS-RO |
The LLM is a security expert analyzing code for a specific vulnerability, with the vulnerability’s definition included. |
| D3 |
FS-RO |
Similar to S6, but includes the definition of the security vulnerability in the system prompt. |
| D4 |
FS-RO |
Like R4, with the addition of the security vulnerability’s definition in the system prompt. |
| D5 |
FS-TO |
Similar to D4, but does not assign a specific role to the LLM in the system prompt. |
Our framework explores four techniques for LLM prompting: (1) zero-shot task-oriented (ZS - TO), (2) zero-shot role-oriented (ZS - RO), (3) few-shot task-oriented (FS - TO), and (4) few-shot role-oriented (FS - RO). Moreover, we divide our set of prompts in the following three categories (all prompts are described in Table III):
@@ 3.4 Datasets
787 |
Out-of-bounds Write |
1 |
C |
79 |
Improper Neutralization of Input During Web Page Generation (‘Cross-site Scripting’) |
2 |
Py |
89 |
Improper Neutralization of Special Elements used in an SQL Command (‘SQL Injection’) |
3 |
Py |
416 |
Use After Free |
4 |
C |
22 |
Improper Limitation of a Pathname to a Restricted Directory (‘Path Traversal’) |
8 |
C |
476 |
NULL Pointer Dereference |
12 |
C |
190 |
Integer Overflow or Wraparound |
14 |
C |
77 |
Improper Neutralization of Special Elements used in a Command (‘Command Injection’) |
16 |
C |
Similar to previous work [9], we create six code scenarios (three pairs of vulnerable and patched scenarios) for each CWE. Moreover, we design our code scenarios with three difficulty levels, (1) easy, (2) medium, (3) hard. Code scenario ‘’ in a specific CWE represents the vulnerable scenario with ‘medium’ difficulty level and ‘’ its patch. The difficulty levels assess how LLMs interact with code of increasing complexity. Easy scenarios consist of simple programs containing only one function and less than 30 lines of code. Medium level scenarios increase the complexity by making the program longer, using different library functions, and adding more than one user input. Hard level introduces scenarios with multiple functions in which functions can be safe on an individual level but when they work together they make the program vulnerable.111Appendix A shows examples of all difficulty levels for ‘CWE-22’.
Real-World CVE Senarios. We leverage a set of real-world Common Vulnerabilities and Exposures (CVEs) from public open source projects to investigate if LLMs are able to identify vulnerabilities in them. Note that existing benchmarks for vulnerability detection [34, 19, 43, 44], cannot be used for this project, as they were released before the cutoff training date of current LLMs, and it is therefore likely that the models saw that data during training. To avoid this potential confounder, we curate 30 code scenarios containing vulnerable and patched versions of 15 CVEs from four open source projects, all published and fixed in 2023, after current LLMs were trained (see Table V).
@@
| gpac |
2023-1452 |
Out-of-Bound Write (787) |
4.5k |
243 |
Mar |
May |
2023-3012 |
NULL Pointer Deref (476) |
2.5k |
398 |
May |
Nov |
2023-23143 |
Out-of-Bound Write (787) |
12.3k |
117 |
Jan |
May |
2023-23144 |
Integer Overflow (190) |
439 |
389 |
Jan |
May |
| libtiff |
2023-2908 |
NULL Pointer Deref (476) |
2.3k |
629 |
Jan |
Nov |
2023-3316 |
NULL Pointer Deref (476) |
159 |
159 |
Jan |
Nov |
2023-26966 |
Out-of-Bound Write (787) |
1.8k |
238 |
Jun |
Nov |
2023-40745 |
Integer Overflow (190) |
2.2k |
757 |
Oct |
Nov |
2023-41175 |
Integer Overflow (190) |
779 |
748 |
Oct |
Nov |
| linux |
2023-40283 |
Use-After-Free (416) |
1.9k |
515 |
Aug |
Nov |
2023-42753 |
Integer Overflow (190) |
628 |
623 |
Sept |
Nov |
2023-42754 |
NULL Pointer Deref (476) |
3.7k |
177 |
Oct |
Nov |
2023-45863 |
Out-of-Bound Write (787) |
1.1k |
565 |
Oct |
Nov |
2023-45871 |
Out-of-Bound Write (787) |
10.1k |
386 |
Oct |
Nov |
| pjsip |
2023-27585 |
Out-of-Bound Write (787) |
784 |
737 |
Mar |
Aug |
@@
Code Augmentations. While standard frameworks exist to evaluate the robustness of LLMs for NLP tasks [45, 46], there is no standard framework to evaluate the robustness of code security related tasks. To fill this gap, we design a set of 150 code scenarios across two categories:
@@
Non-Trivial Augmentations. We design non-trivial code augmentations to perform stress-tests on LLMs to measure their robustness and bias towards semantics of function or variable names, specific library functions, or code security practices. We use combinations of all CWEs defined in Table IV and the six non-trivial code augmentations from Table VI and design 66 code scenarios (12 per NT1-NT4 and 9 per NT5 and NT6) 222See Appendix LABEL:sec:app-robust for details on creation of non-trivial augmented code scenarios.
@@
| T1 |
Rename function parameters randomly |
NT1 |
Change variable names to vulnerability related keywords |
| T2 |
Rename function randomly |
NT2 |
Change the name of a safe function to ‘vulnerable’ function |
| T3 |
Add random unreachable code |
NT3 |
Change the name of an unsafe function to ‘non_vulnerable’ function |
| T4 |
Add random code in comments |
NT4 |
Add a potentially dangerous library function (e.g., ‘strcpy’ or ‘strcat’) but use it in a safe way |
| T5 |
Insert whitespaces |
NT5 |
Use sanitizing functions (e.g., ‘realpath’) in vulnerable code but in a way that it does not resolve the vulnerability |
| T6 |
Add a useless function |
NT6 |
Add hash-defined expressions for safe functions names (e.g., ‘fgets’) but add vulnerable library functions in its body (e.g., ‘gets’) |
| T7 |
Add next-line character |

### v2

# LLMs Cannot Reliably Identify and Reason About Security Vulnerabilities (Yet?): A Comprehensive Evaluation, Framework, and Benchmarks

@@ Abstract
Large Language Models (LLMs) have been suggested for use in automated vulnerability repair, but benchmarks showing they can consistently identify security-related bugs are lacking. We thus develop SecLLMHolmes, a fully automated evaluation framework that performs the most detailed investigation to date on whether LLMs can reliably identify and reason about security-related bugs. We construct a set of 228 code scenarios and analyze eight of the most capable LLMs across eight different investigative dimensions using our framework. Our evaluation shows LLMs provide non-deterministic responses, incorrect and unfaithful reasoning, and perform poorly in real-world scenarios. Most importantly, our findings reveal significant non-robustness in even the most advanced models like ‘PaLM2’ and ‘GPT-4’: by merely changing function or variable names, or by the addition of library functions in the source code, these models can yield incorrect answers in and of cases, respectively. These findings demonstrate that further LLM advances are needed before LLMs can be used as general purpose security assistants.
@@ 1 Introduction
Large Language Models (LLMs), such as OpenAI’s Codex [1], Google’s PaLM2 [2], Meta’s Codellama [3], and StarCoder [4], etc., have demonstrated great potential in performing programming-language related tasks such as code generation, code documentation , and debugging. In 2022, around 1.2 million developers used Copilot, and since then we have witnessed the release of increasingly capable LLM models at a quick pace [2, 5, 4, 3]. LLMs could be particularly useful to help developers with their cybersecurity needs, as humans typically produce and miss many security relevant bugs. This issue was highlighted in the 2022 GitLab Survey [6], noting that “developers do not find enough bugs early enough” and “do not prioritize the bug remediation” when developing. It is pertinent then to investigate if LLMs could be an aid towards early identification of security problems, especially as LLMs have already been suggested for use in automated bug repair [7].
@@
In this paper, we aim to answer the following question: Can LLMs be used as helpful security assistants for vulnerability detection? This is an important question, especially as LLMs are not infallible in security-related tasks, for example introducing vulnerabilities into source code [8, 9] and software testing [10]. Unfortunately, there is no standardized and automated approach to evaluate the performance of LLMs at identifying vulnerable code. We fill this gap by introducing SecLLMHolmes, a generalized, fully automated, and scalable framework to systematically evaluate the performance (i.e., accuracy and reasoning capabilities) of LLMs for vulnerability detection. Our framework tests the capabilities of a given LLM as a security assistant across eight distinct dimensions: (1) deterministic response, (2) performance over range of parameters, (3) diversity of prompts, (4) faithful reasoning, (5) evaluation over variety of vulnerabilities, (6) assessment of various code difficulty levels, (7) robustness to code augmentations, and (8) use in real-world projects.
@@
We apply our framework to eight of the most capable LLMs across 228 code scenarios spanning over 8 most critical vulnerabilities in C and Python, and show that: (a) LLM performance varies widely depending on the model and the prompting technique used, however all models analyzed have a high false positive rate (FPR), and flag code where vulnerabilities have been patched as still vulnerable. (b) the output of LLMs is non-deterministic, with all models changing their answers over multiple runs for one or more of our tests. (c) even when they correctly identify a vulnerability, the reasoning that LLMs provide for this decision is often incorrect, questioning their trustworthiness. (d) LLM chain-of-thought reasoning [11] is not robust, and can be ‘confused’ by even simple code augmentations such as whitespace modification, changing function names, or using different but related library functions. Also, (e) LLMs fail at detecting vulnerabilities in real-world projects. Our study provides significant evidence that LLMs are not yet ready to be used for automated vulnerability detection, and the successful usage of our framework as a benchmark suite by future models would demonstrate meaningful progress in this space.
@@
We develop SecLLMHolmes, a comprehensive framework to test LLMs for their ability to identify and reason about software vulnerabilities. Our framework is fully automated and includes a set of 228 code scenarios, and 17 prompting techniques. We publicly release our framework and dataset111https://github.com/ai4cloudops/SecLLMHolmes, allowing the community to test newly developed LLMs and easily keep track of their progress in being able to identify vulnerabilities.
@@ 2 Background and Related Work
| gpt-4 | GPT-4 | 1.76T | 8,192 | Rem | 09/2021 |
| gpt-3.5-turbo-16k | GPT-3.5 | 175B | 16,385 | Rem | 09/2021 |
| codechat-bison@001 | PaLM2 | 340B | 6,144 | Rem | mid-2021 |
| chat-bison@001 | PaLM2 | 340B | 6,144 | Rem | mid-2021 |
| codellama-7b-instruct | Llama2 | 7B | 100k | Loc | 01/2022 |
| codellama-13b-instruct | Llama2 | 13B | 100k | Loc | 01/2022 |
| codellama-34b-instruct | Llama2 | 34B | 100k | Loc | 01/2022 |
| starchat-beta | StarCoder+ | 15.5B | 8,192 | Loc | 09/2022 |
@@
| Vuln. Code Gen. [8] | 89 | |||||
| Vuln. Code Repair [7] | syn. | |||||
| Limits of ML [30] | res. | |||||
| Transf. Vuln. Det. [34] | res. | |||||
| SecLLMHolmes | 228 |
Recent works have evaluated LLMs for vulnerable code generation [8], code repair [7], and vulnerability detection [34]. However, these approaches are either limited by the number of LLMs, coverage of vulnerabilities, diversity of prompts, range of code complexities, robustness testing, or require manual labor for evaluating LLMs. Most importantly, these studies only evaluate binary responses, checking whether the LLM gives the right label to code snippets (e.g., ‘vulnerable’ or ‘not vulnerable’). In this paper, we present the first comprehensive evaluation framework to evaluate LLMs on the task of vulnerability detection, providing a multi-faceted analysis of the capabilities of LLMs, including going beyond binary decisions and evaluating their reasoning abilities. Table II summarizes evaluation studies of LLMs, highlights their short-comings, and compares against our framework.
In addition to presenting a new state-of-the-art framework, our study uncovers previously unidentified challenges associated with LLMs, including their inability to understand complex code data-flows, the fragility of their COT or step-by-step reasoning to even trivial augmentations, a bias towards security-related functions and variable names that overlooks actual vulnerabilities, a reasoning process that diverges from the methodology of real-world human security experts, particularly in accurately identifying correct root cause of a vulnerability.
3 SecLLMHolmes
Figure 2 presents an overview of our fully automated framework, SecLLMHolmes, which is designed to be applicable to any chat-based LLM. Section 3.1 describes how users can configure an LLM for integration. The core of our framework consists of five pre-defined key components: (i) a set of parameters (Section 3.2), (ii) an extensive set of prompt templates (Section 3.3), (iii) datasets (Section 3.4), and (iv) code augmentations (Section V), all of which facilitate the generation of test prompts. Each prompt is then passed to the LLM to generate a response, and the quality of the response is then analyzed by the (v) ‘Evaluator’ module (Section 3.6).
@@ 3.1 LLM configuration
To scale SecLLMHolmes for any chat-based LLM, we have established three configurable user inputs:
@@ 3.3 Prompt Templates
| S1 | ZS-TO | Code snippet is added to the input prompt with a question about a specific Common Weakness Enumeration (CWE) (e.g., out-of-bound write, path traversal). |
| S2 | ZS-RO | Same as S1, but the LLM is assigned the role of a ‘helpful assistant’. |
| S3 | ZS-RO | Similar to S1, with the LLM acting as a ‘security expert’. |
| S4 | ZS-RO | The LLM is defined as a ‘security expert’ who analyzes a specified security vulnerability, without the question being added to the input prompt. |
| S5 | FS-TO | Similar to S1, but includes a vulnerable example, its patch, and standard reasoning from the same CWE. |
| S6 | FS-RO | Like S4, but also includes a vulnerable example, its patch, and standard reasoning from the same CWE. |
| R1 | ZS-TO | Similar to S1, but begins with ”Lets think step by step” [37] to encourage a methodical approach. |
| R2 | ZS-RO | The LLM plays the role of a security expert with a multi-step approach to vulnerability detection, following a chain-of-thought reasoning. |
| R3 | ZS-TO | A multi-round conversation with the LLM, starting with a code snippet and progressively analyzing sub-components for a security vulnerability like human security-experts. |
| R4 | FS-RO | Similar to S6, but the reasoning for answers involves step-by-step analysis developed by the first author. |
| R5 | FS-RO | Like R2, but includes few-shot examples (from the same CWE) with step-by-step reasoning for detecting vulnerabilities. |
| R6 | FS-TO | Similar to R5, but does not assign a specific role to the LLM in the system prompt. |
| D1 | ZS-TO | Adds the definition of a security vulnerability to the input prompt, followed by a related question. |
| D2 | ZS-RO | The LLM is a security expert analyzing code for a specific vulnerability, with the vulnerability’s definition included. |
| D3 | FS-RO | Similar to S6, but includes the definition of the security vulnerability in the system prompt. |
| D4 | FS-RO | Like R4, with the addition of the security vulnerability’s definition in the system prompt. |
| D5 | FS-TO | Similar to D4, but does not assign a specific role to the LLM in the system prompt. |
SecLLMHolmes explores four techniques for LLM prompting: (1) zero-shot task-oriented (ZS - TO), (2) zero-shot role-oriented (ZS - RO), (3) few-shot task-oriented (FS - TO), and (4) few-shot role-oriented (FS - RO). Moreover, we divide our set of prompts in the following three categories (all prompts are described in Table III):
@@ 3.4 Datasets
| 787 | Out-of-bounds Write | 1 | C |
| 79 | Improper Neutralization of Input During Web Page Generation (‘Cross-site Scripting’) | 2 | Py |
| 89 | Improper Neutralization of Special Elements used in an SQL Command (‘SQL Injection’) | 3 | Py |
| 416 | Use After Free | 4 | C |
| 22 | Improper Limitation of a Pathname to a Restricted Directory (‘Path Traversal’) | 8 | C |
| 476 | NULL Pointer Dereference | 12 | C |
| 190 | Integer Overflow or Wraparound | 14 | C |
| 77 | Improper Neutralization of Special Elements used in a Command (‘Command Injection’) | 16 | C |
Similar to previous work [8], we create six code scenarios (three pairs of vulnerable and patched scenarios) for each CWE. Moreover, we design our code scenarios with three difficulty levels, (1) easy, (2) medium, (3) hard. Code scenario ‘’ in a specific CWE represents the vulnerable scenario with ‘medium’ difficulty level and ‘’ its patch. The difficulty levels assess how LLMs interact with code of increasing complexity. Easy scenarios consist of simple programs containing only one function and less than 30 lines of code. Medium level scenarios increase the complexity by making the program longer, using different library functions, and adding more than one user input. Hard level introduces scenarios with multiple functions in which functions can be safe on an individual level but when they work together they make the program vulnerable.222Appendix A shows examples of all difficulty levels for ‘CWE-22’.
Real-World CVE Senarios. We leverage a set of real-world Common Vulnerabilities and Exposures (CVEs) from public open source projects to investigate if LLMs are able to identify vulnerabilities in them. Note that existing benchmarks for vulnerability detection [33, 18, 42, 43], cannot be used for this project, as they were released before the cut-off training date of current LLMs, and it is therefore likely that the models saw that data during training. To avoid this potential confounder, we curate 30 code scenarios containing vulnerable and patched versions of 15 CVEs from four open source projects, all published and fixed in 2023, after current LLMs were trained (see Table V).
@@
| gpac | 2023-1452 | Out-of-Bound Write (787) | 4.5k | 243 | Mar | May |
| 2023-3012 | NULL Pointer Deref (476) | 2.5k | 398 | May | Nov | |
| 2023-23143 | Out-of-Bound Write (787) | 12.3k | 117 | Jan | May | |
| 2023-23144 | Integer Overflow (190) | 439 | 389 | Jan | May | |
| libtiff | 2023-2908 | NULL Pointer Deref (476) | 2.3k | 629 | Jan | Nov |
| 2023-3316 | NULL Pointer Deref (476) | 159 | 159 | Jan | Nov | |
| 2023-26966 | Out-of-Bound Write (787) | 1.8k | 238 | Jun | Nov | |
| 2023-40745 | Integer Overflow (190) | 2.2k | 757 | Oct | Nov | |
| 2023-41175 | Integer Overflow (190) | 779 | 748 | Oct | Nov | |
| linux | 2023-40283 | Use-After-Free (416) | 1.9k | 515 | Aug | Nov |
| 2023-42753 | Integer Overflow (190) | 628 | 623 | Sept | Nov | |
| 2023-42754 | NULL Pointer Deref (476) | 3.7k | 177 | Oct | Nov | |
| 2023-45863 | Out-of-Bound Write (787) | 1.1k | 565 | Oct | Nov | |
| 2023-45871 | Out-of-Bound Write (787) | 10.1k | 386 | Oct | Nov | |
| pjsip | 2023-27585 | Out-of-Bound Write (787) | 784 | 737 | Mar | Aug |
@@
Code Augmentations. While standard frameworks exist to evaluate the robustness of LLMs for NLP tasks [44, 45], there is no standard framework to evaluate the robustness on code security related tasks. To fill this gap, we design a set of 150 augmented code scenarios, meticulously crafted and reviewed to preserve the ability of human security experts to identify vulnerabilities. These augmentations are organized in two distinct categories:
@@
Non-Trivial Augmentations. We design non-trivial code augmentations to perform stress-tests on LLMs to measure their robustness and bias towards semantics of function or variable names, specific library functions, or code security practices. We use combinations of all CWEs defined in Table IV and the six non-trivial code augmentations from Table VI and design 66 code scenarios (12 per NT1-NT4 and 9 per NT5 and NT6) 333See Appendix LABEL:sec:app-robust for details on creation of non-trivial augmented code scenarios.
@@
| T1 | Rename function parameters randomly | NT1 | Change variable names to vulnerability related keywords |
| T2 | Rename function randomly | NT2 | Change the name of a safe function to ‘vulnerable’ function |
| T3 | Add random unreachable code | NT3 | Change the name of an unsafe function to ‘non_vulnerable’ function |
| T4 | Add random code in comments | NT4 | Add a potentially dangerous library function (e.g., ‘strcpy’ or ‘strcat’) but use it in a safe way |
| T5 | Insert whitespaces | NT5 | Use sanitizing functions (e.g., ‘realpath’) in vulnerable code but in a way that it does not resolve the vulnerability |
| T6 | Add a useless function | NT6 | Add hash-defined expressions for safe functions names (e.g., ‘fgets’) but add vulnerable library functions in its body (e.g., ‘gets’) |
| T7 | Add next-line character | ||